# Gorenstein singularities with $\mathbb{G}_{m}$-action and moduli spaces of holomorphic differentials

### v1

@@ 3.2. $A_{2g+1}$-singularity and $\mathbb{P}\mathcal{H}(g-1,g-1)^{\operatorname{hyp}}$
For $m=2$, $\dim H^{0}(C,(2g-\lambda)(p_{1}+p_{2}))/H^{0}(C,(2g-\lambda-1)(p_{1}+p_{2}))=2$, $1$, or 00, which occurs if $\lambda\leq g-1$, $g\leq\lambda\leq 2g$, or $\lambda>2g$, respectively. We thus obtain that
@@
3.3. $D_{2g+1}$-singularity and $\mathbb{P}\mathcal{H}(0,2g-2)^{\operatorname{hyp}}$
@@
For $m=1$, if $\lambda\geq 2g$, then $\mathcal{F}^{\lambda}H^{0}(C,\omega_{C}^{\log})=0$. Suppose $11\leq\lambda\leq 2g-$ 1. Then $\mathcal{F}^{\lambda}H^{0}(C,\omega_{C}^{\log})=H^{0}(C,(2g-1-\lambda)p_{2})$. Consequently, $\dim H^{0}(C,(2g-1-\lambda)p_{2})/H^{0}(C,(2g-2-\lambda)p_{2})=1$ or 00, where the former occurs if and only if $\lambda=1,3,\ldots,2g-1$. We thus obtain that
@@
For $m=2$, if $\lambda\geq 4g-1$, then $\mathcal{F}^{\lambda}H^{0}(C,2\omega_{C}^{\log})=0$. Next, suppose $22g\leq\lambda\leq 4g-$ 2. Then $\mathcal{F}^{\lambda}H^{0}(C,2\omega_{C}^{\log})=H^{0}(C,(4g-2-\lambda)p_{2})$. Consequently, $\dim H^{0}(C,(4g-2-\lambda)p_{2})/H^{0}(C,(4g-3-\lambda)p_{2})=1$ for $\lambda=2g,2g+2,\ldots,4g-2$. Finally, suppose $11\leq\lambda\leq 2g-$ 1. Then $\mathcal{F}^{\lambda}H^{0}(C,2\omega_{C}^{\log})=H^{0}(C,p_{1}+(4g-2-\lambda)p_{2})$ whose dimension is $3g-\lambda$. Consequently, $\dim H^{0}(C,p_{1}+(4g-2-\lambda)p_{2})/H^{0}(C,p_{1}+(4g-3-\lambda)p_{2})=1$ for all $\lambda$ in this range. We thus obtain that
@@ 3.4. $D_{2g+2}$-singularity and $\mathbb{P}\mathcal{H}(0,g-1,g-1)^{\operatorname{hyp}}$
For $m=2$, if $\lambda\geq 2g+1$, then $\mathcal{F}^{\lambda}H^{0}(C,2\omega_{C}^{\log})=0$. Next, we suppose $g+1\leq\lambda\leq 2g$. Then $\mathcal{F}^{\lambda}H^{0}(C,2\omega_{C}^{\log})=H^{0}(C,(2g-\lambda)(p_{2}+p_{3}))$ which is of dimension $12g-\lambda+$ 1. Consequently, $\dim\mathcal{F}^{\lambda}H^{0}(C,2\omega_{C}^{\log})/\mathcal{F}^{\lambda+1}H^{0}(C,2\omega_{C}^{\log})=1$ for all $\lambda$ in this range. Finally, consider $1\leq\lambda\leq g$. Then $\mathcal{F}^{\lambda}H^{0}(C,2\omega_{C}^{\log})=H^{0}(C,p_{1}+(2g-\lambda)(p_{2}+p_{3}))$ which is of dimension $23g-2\lambda+$ 2. Consequently, $\dim\mathcal{F}^{\lambda}H^{0}(C,2\omega_{C}^{\log})/\mathcal{F}^{\lambda+1}H^{0}(C,2\omega_{C}^{\log})=2$ for all $\lambda$ in this range. We thus obtain that
@@ 3.5. $E_{7}$-singularity and $\mathbb{P}\mathcal{H}(1,3)$
The nontrivial weights are $1$, $2$, and $4$. We thus obtain that
@@
The nontrivial weights are $1$, $2$, $2$, $3$, $4$, $5$, $6$, and $8$. We thus obtain that
@@ 3.6. Elliptic $n$-fold points and $\mathbb{P}\mathcal{H}(0^{n})$
For $n\geq 3$, this singularity is given by the union of $n$ general lines through a point in $\mathbb{A}^{n-1}$. The corresponding stratum is $\mathbb{P}\mathcal{H}(0^{n})$ in genus one, where $0^{n}$ denotes $n$ ordinary marked points, i.e., zeros of order 00.
@@
which is of dimension 00, $1$, or $(m-\lambda)n$ for $\lambda>m$, $\lambda=m$, or $\lambda<m$, respectively.
@@ 3.7. The principal strata
For $m=2$, we have $\dim\mathcal{F}^{\lambda}H^{0}(C,2\omega_{C}^{\log})/\mathcal{F}^{\lambda+1}H^{0}(C,2\omega_{C}^{\log})=2g-2$, $32g-$ 3, $g-1$, or $1$ for $\lambda=1$, $2$, $3$, or $4$, respectively. We thus obtain that
@@
Although all $(Y,q)\in S(1^{2g-2})$ behave the same numerically, $S(1^{2g-2})$ can contain infinitely many non-isomorphic singularity classes. For example, for $g=3$, consider the canonical divisor $p_{1}+p_{2}+p_{3}+p_{4}$ cut out by a line $L$ with a plane quartic curve $C$. Through the construction of the deformation to the normal cone for the divisor at infinity $p_{1}+p_{2}+p_{3}+p_{4}$, the resulting $Y$ consists of four lines joining each $p_{i}$ to a common point $q\in\mathbb{P}^{2}$. Hence, the cross-ratio of $p_{1}$, $p_{2}$, $p_{3}$, and $p_{4}$ in $L\cong\mathbb{P}^{1}$ is an invariant for the singularity $(Y,q)$; see [Bat24, Remark 2.1] for more details. In this case, $\mathbb{P}\mathcal{H}(1^{4})$ admits a slicing over the one-dimensional base of cross-ratios.
@@ 3.8. Toric singularities
#$\#\{(i,j):i,j\geq 0\ {\rm and}\ qi+pj\leq(2k+2-\lambda)b\}.$
If $12k+2-\lambda\geq k+$ 1, namely, $\lambda\leq k+1$, then the dimension of $\mathcal{F}^{\lambda}H^{0}(C,2\omega_{C}^{\log})$ is $1-g+(2k+2-\lambda)b$. From these values, one can also compute $\chi_{2}^{\log}(Y)$ explicitly.
@@ 4. Numerical properties of weights and characters
In this section, we analyze the numerical properties for the weights and characters of the $\mathbb{G}_{m}$-action on $H^{0}(C,m\omega_{C}^{\log})$ for $(C,p_{1},\ldots,p_{n})\in\mathbb{P}\mathcal{H}(\mu)$. Specifically, we will show that the weights for $m=1$ can determine the weights for all $m\geq 2$. Additionally, for later use, we review the relation of the characters of the $\mathbb{G}_{m}$-action and the $\alpha$-invariants of the singularities in the context of the log minimal model program for ¯$\overline{\mathcal{M}}_{g}$.
@@ 4.1. Weights and characters
$N_{m,\lambda}$ $=h^{0}\left(C,m\omega_{C}^{\log}-\sum_{i=1}^{n}l_{\lambda,i}p_{i}\right)-h^{0}\left(C,m\omega_{C}^{\log}-\sum_{i=1}^{n}l_{\lambda+1,i}p_{i}\right).$ (10)
@@
where they are listed with multiplicities and the first $n-1$ weights are 00.
@@ 4.2. The $\alpha$-invariants
$\frac{\chi_{2}^{\log}}{\chi_{1}^{\log}}$ $=\frac{13(1-\alpha)}{2-\alpha}.$ (11)
@@ 5. Singularities of the nonvarying strata
A stratum $\mathbb{P}\mathcal{H}(\mu)$ (or one of its connected components) is called nonvarying, if all Teichmüller curves contained in it have the same slope; equivalently, if they have the same area Siegel–Veech constant, or the same sum of nonnegative Lyapunov exponents. The hyperelliptic strata and a number of strata in low genus are known to be nonvarying; see [CM12, YZ13]. They are also expected to be the only nonvarying strata based on numerical evidences. Additionally, the nonvarying strata possess special properties in affine geometry and intersection theory; see [Che24]. From the viewpoint of this paper, it is natural to speculate that a stratum $\mathbb{P}\mathcal{H}(\mu)$ is nonvarying if and only if $S(\mu)$ consists of a unique isomorphism class $(Y,q)$ of Gorenstein singularities with $\mathbb{G}_{m}$-action. In other words, in this case $\mathbb{P}\mathcal{H}(\mu)$ would be the locus of smooth deformations in the versal deformation space of $(Y,q)$.
In this section, we prove the above speculation for all known nonvarying strata in [CM12, YZ13], as recalled in Theorem 1.2. Additionally, we classify the unique singularity class for each of them, describe their defining equations as well as weighted projective embeddings, and compute their weights, characters, and $\alpha$-invariants. Finally, we prove the second part of Theorem 1.2 in Theorem 5.1.
By the descriptions in Section 3, the conjecture holds for the hyperelliptic strata: $\mathbb{P}\mathcal{H}(2g-2)^{\operatorname{hyp}}$ and $\mathbb{P}\mathcal{H}(g-1,g-1)^{\operatorname{hyp}}$, as well as for the minimal strata up to genus five: $\mathbb{P}\mathcal{H}(4)^{\operatorname{odd}}$, $\mathbb{P}\mathcal{H}(6)^{\operatorname{odd}}$, $\mathbb{P}\mathcal{H}(6)^{\operatorname{even}}$, $\mathbb{P}\mathcal{H}(8)^{\operatorname{odd}}$, and $\mathbb{P}\mathcal{H}(8)^{\operatorname{even}}$, where the Weierstrass semigroup of the unique zero is nonvarying in each case. Therefore, it suffices to consider the remaining nonvarying strata in [CM12, Figures 3 and 4]:
@@
We remark that the conjecture for the nonvarying strata in genus three also follows from the classification in [Bat24, Section 2] (by setting the $\mathbb{G}_{a}$-parameters to be zero in the crimping data in order to have the $\mathbb{G}_{m}$-action well-defined). For completeness, we include the discussion for them as well.
@@
    There are short exact sequences of $\mathbb{C}$-modules:
    $0\to\frac{\widetilde{\mathfrak{m}}^{i}\cap R}{\widetilde{\mathfrak{m}}^{i+1}\cap R}\to\frac{\widetilde{\mathfrak{m}}^{i}}{\widetilde{\mathfrak{m}}^{i+1}}\to(\widetilde{R}/R)_{i}\to 0$
@@
$[\alpha_{1}\ \alpha_{2}\ \alpha_{3}\cdots]$
@@
Moreover, if $\alpha_{i}=0$ and $\alpha_{j}=0$, then $\alpha_{i+j}=0$. We caution the reader not to confuse the singularity gap sequence with the Weierstrass gap sequence of a smooth point.
@@
In the above, (G1) follows from the fact that a Gorenstein curve singularity, except for the node, is indecomposable (see [AFS16, Proposition 2.1]). (G2) follows from the fact that the $\mathbb{G}_{m}$-action on the branch of $t_{i}$ has weight $\ell/(m_{i}+1)$, where $\ell=\operatorname{lcm}(m_{1}+1,\ldots,m_{n}+1)$. The first part of (G3) has been verified in the proof of Theorem 2.6 (iv). The second part of (G3) as well as (G4) and (G5) follow from the fact that the duality residue pairing is perfect, whose germs at the singularity form a graded pairing.
@@ 5.1. The stratum $\mathbb{P}\mathcal{H}(3,1)$
[…]

### v2

@@ 3.2. $A_{2g+1}$-singularity and $\mathbb{P}\mathcal{H}(g-1,g-1)^{\operatorname{hyp}}$
For $m=2$, $\dim H^{0}(C,(2g-\lambda)(p_{1}+p_{2}))/H^{0}(C,(2g-\lambda-1)(p_{1}+p_{2}))=2$, $1$, or 0, which occurs if $\lambda\leq g-1$, $g\leq\lambda\leq 2g$, or $\lambda>2g$, respectively. We thus obtain that
@@
$3.$ $D_{2g+1}$-singularity and $\mathbb{P}\mathcal{H}(0,2g-2)^{\operatorname{hyp}}$
@@
For $m=1$, if $\lambda\geq 2g$, then $\mathcal{F}^{\lambda}H^{0}(C,\omega_{C}^{\log})=0$. Suppose $1\leq\lambda\leq 2g-1$. Then $\mathcal{F}^{\lambda}H^{0}(C,\omega_{C}^{\log})=H^{0}(C,(2g-1-\lambda)p_{2})$. Consequently, $\dim H^{0}(C,(2g-1-\lambda)p_{2})/H^{0}(C,(2g-2-\lambda)p_{2})=1$ or 0, where the former occurs if and only if $\lambda=1,3,\ldots,2g-1$. We thus obtain that
@@
For $m=2$, if $\lambda\geq 4g-1$, then $\mathcal{F}^{\lambda}H^{0}(C,2\omega_{C}^{\log})=0$. Next, suppose $2g\leq\lambda\leq 4g-2$. Then $\mathcal{F}^{\lambda}H^{0}(C,2\omega_{C}^{\log})=H^{0}(C,(4g-2-\lambda)p_{2})$. Consequently, $\dim H^{0}(C,(4g-2-\lambda)p_{2})/H^{0}(C,(4g-3-\lambda)p_{2})=1$ for $\lambda=2g,2g+2,\ldots,4g-2$. Finally, suppose $1\leq\lambda\leq 2g-1$. Then $\mathcal{F}^{\lambda}H^{0}(C,2\omega_{C}^{\log})=H^{0}(C,p_{1}+(4g-2-\lambda)p_{2})$ whose dimension is $3g-\lambda$. Consequently, $\dim H^{0}(C,p_{1}+(4g-2-\lambda)p_{2})/H^{0}(C,p_{1}+(4g-3-\lambda)p_{2})=1$ for all $\lambda$ in this range. We thus obtain that
@@ 3.4. $D_{2g+2}$-singularity and $\mathbb{P}\mathcal{H}(0,g-1,g-1)^{\operatorname{hyp}}$
For $m=2$, if $\lambda\geq 2g+1$, then $\mathcal{F}^{\lambda}H^{0}(C,2\omega_{C}^{\log})=0$. Next, we suppose $g+1\leq\lambda\leq 2g$. Then $\mathcal{F}^{\lambda}H^{0}(C,2\omega_{C}^{\log})=H^{0}(C,(2g-\lambda)(p_{2}+p_{3}))$ which is of dimension $2g-\lambda+1$. Consequently, $\dim\mathcal{F}^{\lambda}H^{0}(C,2\omega_{C}^{\log})/\mathcal{F}^{\lambda+1}H^{0}(C,2\omega_{C}^{\log})=1$ for all $\lambda$ in this range. Finally, consider $1\leq\lambda\leq g$. Then $\mathcal{F}^{\lambda}H^{0}(C,2\omega_{C}^{\log})=H^{0}(C,p_{1}+(2g-\lambda)(p_{2}+p_{3}))$ which is of dimension $3g-2\lambda+2$. Consequently, $\dim\mathcal{F}^{\lambda}H^{0}(C,2\omega_{C}^{\log})/\mathcal{F}^{\lambda+1}H^{0}(C,2\omega_{C}^{\log})=2$ for all $\lambda$ in this range. We thus obtain that
@@ 3.5. $E_{7}$-singularity and $\mathbb{P}\mathcal{H}(1,3)$
The nontrivial weights are $1$, 22, and $4$. We thus obtain that
@@
The nontrivial weights are $1$, 22, 22, 33, 44, 55, 66, and $8$. We thus obtain that
@@ 3.6. Elliptic $n$-fold points and $\mathbb{P}\mathcal{H}(0^{n})$
For $n\geq 3$, this singularity is given by the union of $n$ general lines through a point in $\mathbb{A}^{n-1}$. The corresponding stratum is $\mathbb{P}\mathcal{H}(0^{n})$ in genus one, where $0^{n}$ denotes $n$ ordinary marked points, i.e., zeros of order 0.
@@
which is of dimension 0, 11, or $(m-\lambda)n$ for $\lambda>m$, $\lambda=m$, or $\lambda<m$, respectively.
@@ 3.7. The principal strata
For $m=2$, we have $\dim\mathcal{F}^{\lambda}H^{0}(C,2\omega_{C}^{\log})/\mathcal{F}^{\lambda+1}H^{0}(C,2\omega_{C}^{\log})=2g-2$, 2g32g-3, g1g-1, or $1$ for $\lambda=1$, $2$, 33, or $4$, respectively. We thus obtain that
@@
Although all $(Y,q)\in S(1^{2g-2})$ behave the same numerically, $S(1^{2g-2})$ can contain infinitely many non-isomorphic singularity classes. For example, for $g=3$, consider the canonical divisor $p_{1}+p_{2}+p_{3}+p_{4}$ cut out by a line $L$ with a plane quartic curve $C$. Through the construction of the deformation to the normal cone for the divisor at infinity $p_{1}+p_{2}+p_{3}+p_{4}$, the resulting $Y$ consists of four lines joining each $p_{i}$ to a common point $q\in\mathbb{P}^{2}$. Hence, the cross-ratio of $p_{1}$, $p_{2}$, $p_{3}$, and $p_{4}$ in $L\cong\mathbb{P}^{1}$ is an invariant for the singularity $(Y,q)$; see [Bat24, Remark 2.1] for more details. In this case, $\mathbb{P}\mathcal{H}(1,1,1,1)$ admits a slicing over the one-dimensional base of cross-ratios. We will describe the corresponding singularities and their deformations in detail in Section 5.11.
@@ 3.8. Toric singularities
$\{(i,j):i,j\geq 0\ {\rm and}\ qi+pj\leq(2k+2-\lambda)b\}.$
If $2k+2-\lambda\geq k+1$, namely, $\lambda\leq k+1$, then the dimension of $\mathcal{F}^{\lambda}H^{0}(C,2\omega_{C}^{\log})$ is $1-g+(2k+2-\lambda)b$. From these values, one can also compute $\chi_{2}^{\log}(Y)$ explicitly.
@@ 4. Numerical properties of weights and characters
In this section, we analyze the numerical properties for the weights and characters of the $\mathbb{G}_{m}$-action on $H^{0}(C,m\omega_{C}^{\log})$ for $(C,p_{1},\ldots,p_{n})\in\mathbb{P}\mathcal{H}(\mu)$. Specifically, we will show that the weights for $m=1$ can determine the weights for all $m\geq 2$. Additionally, for later use, we review the relation of the characters of the $\mathbb{G}_{m}$-action and the $\alpha$-invariants of the singularities in the context of the log minimal model program for $\overline{\mathcal{M}}_{g}$.
@@ 4.1. Weights and characters
(10) $N_{m,\lambda}$ $=h^{0}\left(C,m\omega_{C}^{\log}-\sum_{i=1}^{n}l_{\lambda,i}p_{i}\right)-h^{0}\left(C,m\omega_{C}^{\log}-\sum_{i=1}^{n}l_{\lambda+1,i}p_{i}\right).$
@@
where they are listed with multiplicities and the first n1n-1 weights are 0.
@@ 4.2. The $\alpha$-invariants
(11) $\frac{\chi_{2}^{\log}}{\chi_{1}^{\log}}$ $=\frac{13(1-\alpha)}{2-\alpha}.$
@@ 5. Singularities of the nonvarying strata
A stratum $\mathbb{P}\mathcal{H}(\mu)$ (or one of its connected components) is called nonvarying, if all Teichmüller curves contained in it have the same slope; equivalently, if they have the same area Siegel–Veech constant, or the same sum of nonnegative Lyapunov exponents. The hyperelliptic strata and a number of strata in low genus are known to be nonvarying; see [CM12, YZ13]. They are also expected to be the only nonvarying strata based on numerical evidences. Additionally, the nonvarying strata possess special properties in affine geometry and intersection theory; see [Che24]. From the viewpoint of this paper, it is natural to speculate that a stratum $\mathbb{P}\mathcal{H}(\mu)$ is nonvarying if and only if $S(\mu)$ consists of a unique isomorphism class $(Y,q)$ of Gorenstein singularities with $\mathbb{G}_{m}$-action. In other words, in this case $\mathbb{P}\mathcal{H}(\mu)$ would be the locus of smooth deformations in the miniversal deformation space of $(Y,q)$.
In this section, we prove the above speculation for all known nonvarying strata in [CM12, YZ13], as recalled in Theorem 1.2. Additionally, we classify the unique singularity class for each of them, describe their defining equations as well as weighted projective embeddings, and compute their weights, characters, and $\alpha$-invariants. Finally, we prove the remaining parts of Theorem 1.2 in Theorem 5.1 and Theorem 5.5, respectively, as well as prove Theorem 1.3 through Theorem 5.4.
By the descriptions in Section 3, the uniqueness claim holds for the hyperelliptic strata: $\mathbb{P}\mathcal{H}(2g-2)^{\operatorname{hyp}}$ and $\mathbb{P}\mathcal{H}(g-1,g-1)^{\operatorname{hyp}}$, as well as for the minimal strata up to genus five: $\mathbb{P}\mathcal{H}(4)^{\operatorname{odd}}$, $\mathbb{P}\mathcal{H}(6)^{\operatorname{odd}}$, $\mathbb{P}\mathcal{H}(6)^{\operatorname{even}}$, $\mathbb{P}\mathcal{H}(8)^{\operatorname{odd}}$, and $\mathbb{P}\mathcal{H}(8)^{\operatorname{even}}$, where the Weierstrass semigroup of the unique zero is nonvarying in each case. Therefore, it suffices to consider the remaining nonvarying strata in [CM12, Figures 3 and 4]:
@@
We remark that the claim for the nonvarying strata in genus three also follows from the classification in [Bat24, Section 2] (by setting the $\mathbb{G}_{a}$-parameters to be zero in the crimping data in order to have the $\mathbb{G}_{m}$-action well-defined). For completeness, we include the discussion for them as well.
@@
    There are short exact sequences of $\mathbb{C}$-modules: 0 →~mi∩R~mi+1∩R →~mi~mi+1→(~R/R)_i →0.
@@
$\alpha_{1}\ \alpha_{2}\ \alpha_{3}\cdots]$
@@
Moreover, if $\alpha_{i}=0$ and $\alpha_{j}=0$, then $\alpha_{i+j}=0$. We caution the reader not to confuse the gap sequence of a singularity with the Weierstrass gap sequence of a smooth point.
@@
In the above, (G1) follows from the fact that a Gorenstein curve singularity, except for the node, is indecomposable; see [AFS16, Proposition 2.1]. (G2) follows from the fact that the $\mathbb{G}_{m}$-action on the branch of $t_{i}$ has weight $\ell/(m_{i}+1)$, where $\ell=\operatorname{lcm}(m_{1}+1,\ldots,m_{n}+1)$. The first part of (G3) has been verified in the proof of Theorem 2.6 (iv). The second part of (G3) as well as (G4) and (G5) follow from the fact that the duality residue pairing is perfect, whose germs at the singularity form a graded pairing.
@@ 5.1. The stratum $\mathbb{P}\mathcal{H}(3,1)$
In this case, we have $g=3$, $b=2$, and $\delta=4$. According to the preceding observations, the gap sequence is determined by $\alpha_{1}\ \alpha_{2}\ \alpha_{3}\ 1]$, where $\alpha_{4}=1$ is the last nonzero entry and $\sum_{i\geq 1}\alpha_{i}=3$. Since $\alpha_{4}\neq 0$, it implies that $\alpha_{1}\neq 0$ and $\alpha_{2}\neq 0$, and hence the gap sequence is [1 1 0 1][1\ 1\ 0\ 1].